SOME MATLAB FILES

JUHA-MATTI HUUSKO

Abstract. Here are some of my Matlab files.

Contents

1. Image editing  2
1.1. Area calculator  2
1.2. Automatic picture snipping tool  3
2. Complex analysis  4
2.1. Complex map tool 1  4
2.2. Complex map tool 2  5
3. Text editing  6
3.1. Genealogy  6
4. Pictures  6
4.1. Various pictures  6

Most files are here:

http://integraali.com/stuff/

Date: July 27, 2017.

1. Image editing

1.1. Area calculator. Counts pixels in a picture. Given a picture with two domains E, D in different colors and a black background, the function counts the pixels in each domain and calculates the quotient area(E)/ area(D).

Second part of the function assumes that area(D) = π and calculates area(E) by the formula area(E) = π ∗ area(E)/ area(D).

Download in:

https://se.mathworks.com/matlabcentral/fileexchange/62210-area-pix-picture-filename-

1.2. Automatic picture snipping tool. An automatic snipping tool for pictures and formulas. Saves the formula images automatically to .PNG and .EPS format.

If a pdf document is translated through OCR-software and Google Translate into a Word/L^{A}TEXfile, the formulas can be added in a simple way.

Download in:

https://se.mathworks.com/matlabcentral/fileexchange/62211-formula-snip-filename-filetype-opt-

2. Complex analysis

2.1. Complex map tool 1. Maps dots by a complex map z ↦ f(z). Dots are given by graphical input. In the example picture, the map z ↦ z^{2} is applied.

2.2. Complex map tool 2. Maps an polygonal chain by a complex map z ↦ f(z). Vertices are given by graphical input. In the example picture, the map z ↦ e^{z} is applied.

3. Text editing

3.1. Genealogy. I have written some genealogy details to a file. The function showgenealogy.mat shows the information nicely. The graphics need to be improved.

Download in:

http://integraali.com/stuff/showgenealogy.m

4. Pictures

4.1. Various pictures. During my academics, I have made the following pictures:

Department of Physics and Mathematics, University of Eastern Finland, P.O. Box 111, FI-80101 Joensuu, Finland

E-mail address: juha-matti.huusko@uef.fi